• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name W. Korsch

    Log entry time 12:53:06 on August 5,2001

    Entry number 69421

    keyword=Late entry: Network/DAQ problems


    After the network problems this morning the software person (CC) on call
    (Mike Memory ??) came and checked out the network. According to his data logs the
    cause of the network problem was due to several (self) reboots of the network switches hallasc-cat and halla-cat located in the Hall A electronic room (middle
    room of counting house). It turned out that both switches were connected to
    the same UPS and so the expert suggested to remove the UPS (nothing else
    was connected to it) and use the power directly from an outlet (he promised to
    bring a new UPS tomorrow - we should make sure that it'll happen).
    After the reset of the switches the EPICS readout of the right arm spectrometer
    still didn't work. The accelerator software person on call was very difficult
    to reach (>1 hour response time), but finally showed up around 8:30 am.
    Supposedly he checked everything and came to the conclusion that we
    should contact J. Gomez (~9:45am).

    We also noticed that the right arm DAQ did not collect any data (rates etc
    were showing 0 al the time after a run started, left arm was ok). Contacted
    Bob Michaels several times and he had suggested earlier already to reboot
    the VME crates in the right arm.
    Requested access around 10:00 am and reset the VME crate on the lower level
    of the detector hut (162, slow controls). EPCIS readout worked again,
    but still problems with right arm DAQ (not rates...). Contacted
    Bob Michaels again and he started looking into it from home.
    The problem was related to the rcServer process. Bob rebooted adaql2
    from home and then we started rcServer and coda again.
    Took a short pulser run in the right arm (run# 21609, ps8=10) and
    everything seemed ok (see entry 69403 by Bob Michaels).
    Resumed data taking aroun 11:40 am.